i f a,b are the roots of a tan 2 theta + b sec 2 theta =c prove that tan a .tan b =b-c/b+c

i
f a,b are the roots of a tan 2 theta + b sec 2 theta =c prove that tan a .tan b =b-c/b+c
 

Grade:11

1 Answers

Utsav Basu
6 years ago
α and β are roots of atan2 θ + bsec2 θ – c = 0
Now,
     atan2 θ + bsec2 θ – c = 0
or, atan2 θ + b(1 + tan2 θ) – c = 0
or, (a + b) tan2 θ + (b – c) = 0
 
Thus, tan α * tan β = (b – c)/(a + b)
 
I am sorry, but this is the correct answer. I suppose your question is wrong.